### v1

# Predictive Analysis for Optimizing Port Operations

Abstract
Maritime transport is a pivotal logistics mode for the long-distance and bulk transportation of goods. However, the intricate planning involved in this mode is often hindered by uncertainties, including weather conditions, cargo diversity, and port dynamics, leading to increased costs. Consequently, accurately estimating vessel total (stay) time at port and potential delays becomes imperative for effective planning and scheduling in port operations. This study aims to develop a port operation solution with competitive prediction and classification capabilities for estimating vessel Total and Delay times. This research addresses a significant gap in port analysis models for vessel Stay and Delay times, offering a valuable contribution to the field of maritime logistics. The proposed solution is designed to assist decision-making in port environments and predict service delays. This is demonstrated through a case study on Brazil ports. Additionally, feature analysis is used to understand the key factors impacting maritime logistics, enhancing the overall understanding of the complexities involved in port operations.
Keywords: Supervised Learning, Machine Learning, Feature Importance, Supply Chain, Port Analysis.
1 Introduction
In the intricate landscape of global maritime logistics, an often overlooked yet substantial segment of a vessel’s journey unfolds within the confines of port facilities. While traditional narratives tend to focus on vessels (ships) navigating the open waters [Ogura et al., 2021, Kolley et al., 2022, Mekkaoui et al., 2022], a substantial portion of their journey involves intricate maneuvers, cargo handling, and operational intricacies within ports [Nguyen et al., 2019]. Surprisingly, this dualistic nature of a vessel’s journey—both at sea and within port—has received scant attention in the existing literature. The temporal dynamics, efficiency challenges, and optimization opportunities inherent in the port-centric leg of a vessel’s journey warrant closer scrutiny [Abreu et al., 2022b]. This gap in scholarly exploration calls for an in-depth examination, shedding light on the pivotal role that port activities play in the broader context of maritime transportation, supply chain efficiency, and global trade networks [Dimitrios et al., 2023].
The optimization of port operations holds critical importance in the context of globalization, where nations strive to maintain their competitiveness and economic resilience in an interconnected world. Efficient port operations are essential not only for facilitating trade and commerce but also for accommodating the growing flow of goods and passengers associated with global tourism and cultural exchanges. The integration of marine traffic information through APIs has emerged as a key technological advancement, offering real-time insights into vessel movements and enabling proactive decision-making to enhance port efficiency on a global scale.
While the specific focus of this project is on Brazil ports, the broader implications extend to a global context. Ports worldwide face similar challenges related to the optimization of operations, the impact of globalization on trade flows, and the integration of advanced technologies for enhanced efficiency [Çağatay Iris and Lam, 2019, Izaguirre et al., 2020a]. By studying the case of Brazil, with its rich maritime history and strategic positioning in global trade, valuable insights can be gained that are applicable to port operations and maritime logistics across diverse regions and economies.
The main contribution of the paper is as follows:
  • We define the problem for port operational analysis which is lacking in the current literature.
  • We execute an exhaustive comparison of methods with multiple metrics for prediction and classification of Total time (stay time) and Delay time at port.
  • We discover key factors that affect port operations using feature importance.
The paper is structured as follows: Section 2 provides a comprehensive review of the literature relevant to the study. Section 3 discusses the Brazil port case study and outlines the proposed methods used for port analysis. Section 4 dives into the results and share keys features that impact the port operations. We finish with Section 5 that offers the conclusions from the study, and suggests avenues for future research.
2 Literature Review
The Berth Allocation Problem (BAP) is a critical challenge in port operations, aiming to optimize the allocation of berthing facilities for the incoming ships while considering various restrictions and objectives [Bierwirth and Meisel, 2015]. The estimation of different durations for the vessels is a crucial factor in BAP mathematical formulations, as it directly influences berth allocation decisions. Therefore, the development of efficient prediction and classification models for Total time is strategically important for addressing the BAP. Studies [Guo et al., 2023, Rodrigues and Agra, 2022] have explored novel algorithms and optimization techniques to address the BAP, emphasizing the importance of accurate Total time predictions in optimizing berth allocations.
Predictive Model Scheduling (MPS) is another area of interest in port operations, focusing on the development of predictive models to optimize scheduling processes for port activities [Dulebenets, 2018, van Boetzelaer, 2013]. By leveraging predictive analytics, MPS aims to enhance the efficiency and effectiveness of port operations by forecasting demand, optimizing resource allocation, and improving overall performance [Weerasinghe et al., 2023]. These works have explored the application of machine learning algorithms for predictive scheduling in container terminals, demonstrating the potential of these methods in improving port productivity.
Performance management of ports and vessels is a critical aspect of maritime logistics, involving the monitoring, analysis, and optimization of key performance indicators (KPIs) to ensure operational efficiency and competitiveness. [Farag and Ölçer, 2020, López-Bermúdez et al., 2019] emphasizes the importance of performance management systems for ports, highlighting the need for real-time data analytics and decision support tools to enhance operational performance. Regarding port and ship characteristics related to faults and accidents, the literature has focused on risk assessment and mitigation strategies to improve safety and reliability in port operations [Wang et al., 2021, Cao et al., 2023].
In addition to the aforementioned areas, recent literature has also emphasized the importance of sustainability and green practices in port operations [Izaguirre et al., 2020b, Lim et al., 2019]. With increasing global awareness of environmental issues, there has been a growing focus on developing sustainable and eco-friendly port operations. Sustainable port initiatives aim to minimize the environmental impact of port activities, reduce carbon emissions, and promote the use of renewable energy sources.
Despite the significant work on port operations, there is a absence of studies on models for predicting and classifying Total time and Delay time of vessels at ports, which highlights the need for research in this area. By addressing this gap, the proposed study aims to contribute valuable insights that can inform decision-making processes in port management, ultimately improving the efficiency and effectiveness of port operations. This information will be crucial for solving other port related problems such as the BAP, as accurate predictions of Total time and Delay time can enhance the applicability of BAP solutions in real-world port environments. Moreover, these insights can support management decisions related to the planning, allocation, and scheduling of loading and unloading operations, contributing to overall operational excellence in port logistics.
3 Case Study and Methodology
The Case study of Brazil port is from [Abreu et al., 2022b, Abreu et al., 2022a]. Brazil’s marine logistics and port operations are crucial components of its economy due to the country’s extensive coastline and reliance on maritime transport for international trade. Brazil boasts one of the world’s most extensive coastlines, stretching over 7,400 kilometers, with numerous ports facilitating the import and export of goods. The country’s ports play a pivotal role in connecting Brazil to global markets, serving as crucial gateways for international trade [Cabral and Ramos, 2018, López-Bermúdez et al., 2019]. However, port logistics in Brazil are complex due to inflexible regulations and the need for extensive information, impacting delivery times and costs [da Veiga Lima and de Souza, 2022, Galvao et al., 2017]. Maritime logistics and port operations are vital for Brazil’s economy for several reasons: Trade Facilitation, Economic Growth, Job Creation, Transportation Network, and Strategic Importance.
Overall, the efficient functioning of Brazil’s marine logistics and port operations is essential for the country’s economic prosperity, trade competitiveness, and integration into the global economy [Costa et al., 2022, Cacho et al., 2021]. It represents a critical component of Brazil’s infrastructure and contributes significantly to its overall development.
We have translated the data from Portuguese to English. The data encompasses a wide array of characteristics related to maritime cargo transportation and port operations from a period of 11/2017-12/2018. They are categorized into four main groups: cargo characteristics, geographic characteristics, operation characteristics, and stay (Total) time characteristics.
Starting with cargo characteristics, the dataset includes details such as the type of cargo (ST Nature Cargo), the quantity of cargo moved in Twenty-foot Equivalent Units (TEU), the net weight of containerized cargo (VL Containerized Load Weight), and the overall weight of transported cargo in tons (Moved Value). These variables provide essential insights into the nature and volume of cargo being transported, which are crucial for understanding the logistical demands of port operations.
Moving on to geographic characteristics, the dataset encompasses information about the source and destination of the cargo, the classification of inland transportation routes, the specific mooring berth utilized, the name of the mooring port and port complex, the city and province where the port is located, and the geographical region (such as Southeast, Northeast, North, or South). These variables offer geographic context and detail the locations and routes involved in the transportation process, which can impact logistical planning and operational decision-making.
The operation characteristics section includes variables that detail various aspects of the cargo operations themselves. This includes the classification of cargo operations, the status of containers (whether they are full, empty, or carrying loose cargo), indicators for cabotage and long course operations, the sense of the operation (whether it is disembarked or embarked), the type of port authority (public or private), the classification of mooring navigation, and the nationality of the shipowner (whether Brazilian or foreign). These variables provide insights into the operational dynamics and regulatory aspects of cargo transportation.
Finally, the stay (Total) time characteristics section includes variables related to the duration of a vessel’s stay in port. This encompasses details such as the arrival date and hour of the ship, the mooring date and hour, the unblocking date and hour, the start and end dates and hours of operations, the month of mooring, and various time durations (T1-Wait time for the moat, T2-Wait time for start of operation, T3-Operation time, T4-Wait time for unberthing, T2+T3+T4-Mooring time, and T1+T2+T3+T4-Total time) that reflect different aspects of the vessel’s stay in port. These variables are critical for understanding the temporal aspects of port operations, including the time spent waiting, in operation, and overall stay duration. We define Delay time as Total time-Operation time (T1+T2+T4).
We define the Total time (stay time) class using Fig. 1, similar to [Abreu et al., 2022b]. Class 0 (Low) is defined for vessels with Total time at port for 0-25 hours, Class 1 (Medium) is defined for 25-50 hours, Class 2 (High) includes 50-75 hours and Class 3 (Very High) is $>$75 hours. For Delay Class, we use the plots as well as summary statistics (mean Delay $\approx$ 10 hours, variance of Delay $\approx$125), and define class 0 as Delay time $<$ 24 hours and Class 1 as Delay time $>$ 24 hours.
Figure 1: Distribution of Total time and Delay time at port
In the realm of supervised learning for prediction and classification, tree-based methods and statistics-based methods represent two prominent approaches, each with its strengths and applications. Tree-based algorithms, such as Random Forest, Extreme Gradient Boosting (XGBoost), AdaBoost, Light Gradient Boosting Machine (LightGBM), and Extra Trees, are known for their ability to handle complex relationships in data and provide robust predictions. Random Forest, for instance, constructs multiple decision trees and aggregates their outputs to improve accuracy and reduce overfitting [Breiman, 2001]. XGBoost, a gradient boosting algorithm, excels in optimizing complex objective functions and has been widely used in competitions and industry applications for its performance [Chen and Guestrin, 2016]. AdaBoost focuses on iteratively boosting the performance of weak learners to achieve strong predictive models [Freund and Schapire, 1997]. LightGBM is known for its efficiency in handling large datasets and has become popular for its speed and accuracy [Ke et al., 2017]. Extra Trees, or Extremely Randomized Trees, is an ensemble learning method that builds multiple decision trees and averages their predictions, providing robustness against noise and overfitting [Geurts et al., 2006]. Additionally, feature importance analysis in tree-based models helps identify the most influential variables in the prediction process, providing insights into the underlying patterns in the data [Louppe et al., 2013].
On the other hand, statistics-based methods like regression, Support Vector Machines (SVM), Bayesian classifiers, k-Nearest Neighbors (kNN), and Elastic Net offer a different set of tools for prediction and classification tasks. Regression techniques, including linear regression and its variants, are widely used for predicting continuous outcomes based on the relationships between variables [James et al., 2013]. SVM, a powerful classification method, aims to find the optimal hyperplane that best separates data points into different classes in a high-dimensional space [Cortes and Vapnik, 1995]. Bayesian classifiers, based on Bayes’ theorem, are probabilistic models that calculate the probability of a sample belonging to a particular class [Hand and Yu, 2001]. kNN, a simple yet effective algorithm, classifies a data point based on the majority class of its nearest neighbors in the feature space [Altman, 1992]. Elastic Net, a regularization technique combining L1 (Lasso) and L2 (Ridge) penalties, is particularly useful for feature selection and dealing with multicollinearity in regression problems [Zou and Hastie, 2005]. Each of these statistics-based methods has its own assumptions, strengths, and limitations, making them suitable for different types of datasets and prediction tasks.
In both prediction and classification tasks, various metrics are used to evaluate the performance of machine learning models. For prediction tasks, common metrics include Root Mean Squared Error (RMSE), Mean Absolute Error (MAE), Mean Squared Error (MSE), R-squared ($R^{2}$), Root Mean Squared Logarithmic Error (RMSLE), and Mean Absolute Percentage Error (MAPE). MAE measure the average magnitude of the errors between predicted and actual values. MSE is the average of the squared differences between predicted and actual values (RMSE is square root of MSE). $R^{2}$ represents the proportion of the variance in the dependent variable that is predictable from the independent variables. RMSLE is particularly useful for tasks where the target variable spans several orders of magnitude. MAPE measures the percentage difference between predicted and actual values. For all the metrics, lower values indicate better fit whereas for $R^{2}$ values close to 1 are considered good.
For classification tasks, metrics such as Accuracy, Precision, Recall, Area Under the Receiver Operating Characteristic Curve (AUC), F1 score, and Cohen’s Kappa are commonly used. All this values can be derived from the confusion matrix. Accuracy measures the proportion of correctly classified instances over the total number of instances. Precision measures the proportion of true positive predictions among all positive predictions, focusing on the accuracy of positive predictions. Recall (also known as sensitivity) measures the proportion of true positive predictions among all actual positives, focusing on the model’s ability to detect positive instances. AUC measures the model’s ability to distinguish between classes. F1 score is the harmonic mean of precision and recall, providing a balance between the two metrics. Cohen’s Kappa measures the agreement between predicted and actual classifications, accounting for the possibility of the agreement occurring by chance. For all the metrics, higher values indicate good classification performance.
We use Cross-validation [Kohavi, 1995] to assess the model’s performance and tune its hyperparameters by repeatedly splitting the dataset into subsets for training and validation. It helps prevent overfitting by evaluating the model’s performance on multiple subsets of the data, allowing for more reliable hyperparameter tuning and regularization.
In the context of port operations, the aforementioned machine learning methods are applied to predict and classify Total time and Delay time, which are critical factors in port logistics. By leveraging these techniques, we gain valuable insights into the complex dynamics of port operations.
4 Results
We split the data (n=153331) into 80% (122664) for training and 20% (30667) for testing. We consider only 24 features and remove any feature information related to calculation of Total time and Delay time at the port. To ensure we don’t over fit and to pick the best hyperparameters, we perform 10-fold cross validation (CV). We first look into the prediction of Total time and Delay time, followed by classification of Total time and Delay time. The experiments were performed on an 8-core M2 CPU.
Also, we perform feature analysis for a comprehensive understanding of the key factors influencing port operations. Addressing these factors is crucial for optimizing the Berth Allocation Problem (BAP), Predictive Model Scheduling (MPS), port performance management, and risk assessment. By enhancing these aspects, ports can achieve greater sustainability, aligning with the goal of promoting green practices.
Table 1 shows the performance of different Machine Learning approaches for Total time prediction. We see a general trend that tree based methods are performing better than statistical approaches on this task. This might be because of the non linear nature in the data. We observe that Random Forest (RF) performs the best, has the lowest RMSE (10.4377) and also has superior values in other metrics in comparison to rest of the methods. Performance on the test set for RF is similar, where the RMSE is 10.2577 (other metrics are also in close range). This is visible from Fig. 3, where the predicted and true values are near to the identity line with $R^{2}$ value of 0.8287.
Fig. 2 shows the important features for the Total time prediction, and we can observe that berth feature plays the most crucial role due to some berths being more busy and less efficient compared to others, followed by other features like moved value, VL Containerized Load Weight, type of mooring navigation, source, destination, month, etc. Berths play a vital role in port operations and hence, efficient berth management is essential for optimizing port capacity, throughput, and overall operational efficiency.
Model MAE MSE RMSE $R^{2}$ RMSLE MAPE
Random Forest 5.3336 108.9832 10.4377 0.8241 0.2992 0.1358
Extra Trees 5.0287 121.0471 10.9998 0.8046 0.3112 0.1282
Extreme Gradient Boosting 7.4763 125.7988 11.2152 0.7969 0.3343 0.1974
Light Gradient Boosting Machine 8.3872 147.0024 12.1234 0.7627 0.3648 0.2256
Decision Tree 5.4081 165.8789 12.8762 0.7322 0.3469 0.1378
K Nearest Neighbors 8.1328 184.7665 13.5920 0.7017 0.3968 0.2074
Gradient Boosting 10.4030 211.0671 14.5265 0.6593 0.4238 0.2728
Bayesian Ridge 11.8099 273.8831 16.5479 0.5579 0.4772 0.3203
Ridge Regression 11.8092 273.8808 16.5479 0.5579 0.4773 0.3202
Linear Regression 11.8089 273.8825 16.5479 0.5579 0.4774 0.3201
Elastic Net 12.8900 314.4793 17.7318 0.4924 0.5022 0.3847
Lasso Regression 12.9111 315.2336 17.7531 0.4911 0.5023 0.3845
Huber Regressor 12.5049 315.4952 17.7600 0.4907 0.5316 0.3412
AdaBoost 18.0348 502.1311 22.3968 0.1889 0.6748 0.7616
Table 1: Comparing 10-fold CV results of different methods with multiple metrics (sorted using RMSE) for Total time prediction at port
Figure 2: Feature importance for of Total time prediction at port
Figure 3: R-squared plot on test data for Total time prediction at port
We see a similar story from Table 2 where the tree based methods are performing better for predicting the Delay time at port. Random Forest (RF) again performs the best, with an RMSE of 4.2179. The test set performance for RF gives a RMSE of 4.3707 (other metrics are also in similar range). From Fig. 3, we can see that the predicted and true values are not close to the identity line ($R^{2}$ is 0.6086). The prediction of Delay time might be limited by the feature set, as it is known that weather information, and other external factors play a key role with respect to delay.
Fig. 4 shows the important features for the Delay time prediction, and we can observe that moved value, VL Containerized Load Weight, and Berth are key features followed by other features like source, destination, month, etc. It is intuitive that vessels with higher moved value, and VL Containerized Load Weight goes through longer operations leading to build up in Delay time.
Model MAE MSE RMSE $R^{2}$ RMSLE MAPE
Random Forest 0.8102 17.8800 4.2179 0.6160 0.4997 0.5579
Extreme Gradient Boosting 1.2285 18.1063 4.2470 0.6112 0.6103 0.7758
Extra Trees 0.7304 19.8816 4.4488 0.5729 0.4981 0.5186
Light Gradient Boosting Machine 1.3226 20.9450 4.5700 0.5502 0.6241 0.8042
Decision Tree 0.7571 26.6984 5.1538 0.4255 0.5464 0.5569
Gradient Boosting 1.7824 29.9394 5.4656 0.3574 0.7811 0.9428
K Nearest Neighbors 1.3783 35.7453 5.9735 0.2310 0.6818 0.7387
Bayesian Ridge 2.3272 42.8552 6.5377 0.0815 0.9559 1.1375
Ridge Regression 2.3273 42.8554 6.5378 0.0815 0.9564 1.1423
Linear Regression 2.3271 42.8556 6.5378 0.0815 0.9564 1.1425
Elastic Net 2.0899 44.3601 6.6513 0.0494 0.8636 0.9716
Lasso Regression 2.1192 44.7360 6.6794 0.0414 0.8685 0.9946
Huber Regression 1.2250 48.0613 6.9232 -0.0299 0.7032 0.9955
AdaBoost 3.2974 54.0987 7.3534 -0.1708 1.1535 1.6455
Table 2: Comparing 10-fold CV results of different methods with multiple metrics (sorted using RMSE) for Delay time prediction at port
Figure 4: Feature importance for of Delay time prediction at port
Figure 5: R-squared plot on test data for Delay time prediction at port
In the classification task, the performance of various Machine Learning methods for (Multi-Class) classification of Total time is depicted in Table 3. A consistent trend is observed where tree-based methods outperform statistical approaches, potentially due to the complex nature of the data. Specifically, Extreme Gradient Boosting (XGBoost) demonstrates the best performance, achieving the highest Accuracy (0.8424) and superior values across other metrics compared to alternative methods. This is 10% higher than the result in [Abreu et al., 2022b]. The test set results further confirm XGBoost’s efficacy, with an Accuracy of 0.8415 and other metrics exhibiting similarly favorable performance as derived from the confusion matrix in Fig. 7. Also, we observe that is it easier to detect the outer classes, Class 0 (Low) and Class 3 (Very High), compared to the other two.
Additionally, Fig. 6 showcases the key features influencing Total time classifcation. Notably, the geographical region (north), Port Complex (Manaus), Type of Mooring Navigation (inside) emerges as the most influential features, followed by factors such as City, Mooring port, Inland transportation route, type of mooring navigation (Long Haul), Berth, etc. The significance of the geographical features is related to the huge coastline of Brazil and the complex dynamics of port operations.
Model Accuracy AUC Recall Precision F1 Kappa
Extreme Gradient Boosting 0.8424 0.9716 0.8424 0.8426 0.8424 0.7797
Extra Trees 0.8386 0.9453 0.8386 0.8386 0.8386 0.7743
Random Forest 0.8360 0.9634 0.8360 0.8361 0.8360 0.7707
Decision Tree 0.8220 0.8775 0.8220 0.8219 0.8219 0.7511
Light Gradient Boosting Machine 0.7978 0.9518 0.7978 0.7981 0.7977 0.7165
Gradient Boosting 0.7210 0.9051 0.7210 0.7235 0.7190 0.6046
K Nearest Neighbors 0.7160 0.8962 0.7160 0.7171 0.7156 0.6009
Linear Discriminant Analysis 0.6344 0.8353 0.6344 0.6355 0.6281 0.4788
Ridge Classifier 0.6296 0.8194 0.6296 0.6355 0.6117 0.4661
Logistic Regression 0.5997 0.8135 0.5997 0.6183 0.5626 0.4212
Ada Boost 0.5511 0.7245 0.5511 0.5447 0.5457 0.3687
SVM 0.5412 0.7749 0.5412 0.5808 0.5159 0.3382
Naive Bayes 0.5342 0.7675 0.5342 0.5769 0.5136 0.3538
Quadratic Discriminant Analysis 0.3429 0.7174 0.3429 0.4488 0.2962 0.1611
Table 3: Comparing 10-fold CV results of different methods with multiple metrics (sorted using Accuracy) for Total time classification at port
Figure 6: Feature importance for of Total time classification at port
Figure 7: Confusion matrix on test data for Total time classification at port
Table 4 presents the binary classification performance on Delay time. The tree-based methods are again outperforming statistical approaches. XGBoost gives the best performance by achieving the highest Accuracy (0.9511) and best values across other metrics compared to other methods. XGBoost achieves an Accuracy of 0.9505 on the test set and other metrics showing close performance, as illustrated by the confusion matrix in Fig. 9.
Furthermore, Fig. 8 highlights the key features influencing Delay time classification, factors such as Flag Long Course, Flag Cabotage, Owner’s Nationality and type of mooring navigation (Long Haul) emerge as the most influential. This is followed by variables like Port Complex, Load Operation Type, ST Cargo Nature, and Month. Features related to operation characteristics are having the most impact on Delay class.
Model Accuracy AUC Recall Precision F1 Kappa
Extreme Gradient Boosting 0.9511 0.9798 0.6923 0.7507 0.7202 0.6935
Extra Trees 0.9498 0.9365 0.7014 0.7348 0.7176 0.6901
Random Forest 0.9491 0.9701 0.6853 0.7364 0.7098 0.6820
Decision Tree 0.9459 0.8394 0.6907 0.7079 0.6991 0.6694
Light Gradient Boosting Machine 0.9389 0.9688 0.4375 0.7996 0.5652 0.5354
K Nearest Neighbors 0.9228 0.8741 0.4426 0.6027 0.5104 0.4695
Gradient Boosting 0.9166 0.8990 0.1295 0.7371 0.2196 0.1979
Ridge Classifier 0.9097 0.8418 0.0360 0.5570 0.0676 0.0573
Logistic Regression 0.9079 0.8080 0.0436 0.4363 0.0792 0.0638
Linear Discriminant Analysis 0.9059 0.8159 0.1250 0.4385 0.1945 0.1607
Ada Boost 0.9056 0.8277 0.0684 0.3893 0.1162 0.0917
SVM 0.8135 0.8375 0.2273 0.3695 0.1117 0.0688
Naive Bayes 0.5680 0.7600 0.8872 0.1606 0.2720 0.1394
Quadratic Discriminant Analysis 0.1167 0.5137 0.9978 0.0932 0.1705 0.0049
Table 4: Comparing 10-fold CV results of different methods with multiple metrics (sorted using Accuracy) for Delay time classification at port
Figure 8: Feature importance for of Delay time classification at port
Figure 9: Confusion matrix on test data for Delay time classification at port
5 Conclusion
In conclusion, the development of a comprehensive prediction and classification solution for estimating vessel stay and Delay times in maritime logistics represents a significant step forward in addressing the complexities of port operations. By providing tools to accurately predict and classify vessel behavior, this research contributes to more efficient and effective decision-making in port management. The wide-ranging applications of the proposed solution, from ship scheduling to service delay predictions to risk assessment, demonstrate its potential to enhance overall port performance and optimize supply chain operations. Additionally, the incorporation of feature analysis enriches our understanding of the factors influencing maritime logistics, where predictions are influenced by Berth and Cargo Characteristics features and classification tasks are impacted by Geographical features (Total time) and operation characteristic features (Delay time). This offers insights that can be leveraged to further improve the efficiency of port operations and make them more resilient. Since Berth plays a key role in prediction of Total time and Delay time, using a Top-Bottom approach for developing individual model for ports and then berths would be beneficial in understanding how these ports and berths can be made more efficient.
Focus on refining the prediction and classification models to account for additional variables and complexities in maritime logistics is a key area of future work. This could involve integrating real-time data streams, such as weather forecasts, vessel tracking, other vessel information, to enhance the accuracy and responsiveness of the models. Furthermore, exploring the application of more advanced techniques, such as deep learning and reinforcement learning, may offer new avenues for improving the predictive capabilities of the models. Additionally, efforts to expand the scope of the research beyond individual ports to encompass broader regional or global maritime networks could provide valuable insights into optimizing port operations within larger logistical contexts. Overall, continued research in this area has the potential to drive significant advancements in maritime logistics, leading to more efficient and resilient port operations in the future.